1. What is the heat (q) needed by water with molar heat of vaporization of 40.7 kJ if it is 50.0 g in mass and boils at 100°C? Note that the molar mass of water is 18.0 g/mol.

2. How will you compare it with the heat needed by alcohol with molar heat of vaporization 44.0 kJ at 35.0 g and boils at 82.5°C? Note that the molar mass of isopropyl alcohol is 60.1 g/mol.


1
Expert's answer
2021-04-19T04:34:00-0400

1. q = "\Delta H_{vap} ×" "\frac{mass}{molar mass}"


q = 40.7 × "\frac{50}{18}"

q = 113.05 KJ


2. q = "\Delta H_{vap} ×" "\frac{mass}{molar mass}"

q = 44 × "\frac{35}{60.1}"


q = 25.62 KJ